7.00×10−3 mol of HBr are dissolved in water to make 15.0 L of solution. What is the concentration of hydroxide ions, [OH−], in this solution?
Answer #1

[HBr] = mol of HBr / volume in L
= 7.00*10^-3 mol / 15.0 L
= 4.67*10^-4 M

So,
[H+] = 4.67*10^-4 M

Given:
[H+] = 4.67*10^-4 M

use:
[OH-] = Kw/[H+]
Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC
[OH-] = (1.0*10^-14)/[H+]
[OH-] = (1.0*10^-14)/(4.67*10^-4)
[OH-] = 2.141*10^-11 M
Answer: 2.14*10^-11 M